Using Charpit’s method, find a complete integral of the following equations: xpq + yq^2 = 1


1
Expert's answer
2021-01-19T12:32:41-0500

To find a complete integral of "f=xpq+yq^2-1=0"

"\implies" The auxiliary equations are "\frac{dx}{xq}=\frac{dy}{2qy}=\frac{dz}{xpq+2yq^2}=\frac{dp}{-pq}=\frac{dq}{-q^2}". Hence "\frac{dp}{-pq}=\frac{dq}{-q^2}" gives p-aq. Thus g=p-aq=0. Solving f=0 and g=0 for p and q, we get "q=\frac{1}{\sqrt{ax+y}}, p=\frac{q}{\sqrt{ax+y}}". Hence "dz=\frac{adx+dy}{\sqrt{ax+y}}". Thus "(z+b^2)=4(ax+y)" is the required complete integral.